Magnesium metal reacts with hydrochloric acid to produce magnesium chloride and hydrogen gas.  If 34.0 grams of hydrochloric acid reacts, how many grams of hydrogen gas is produced?



9.0 moles of magnesium phosphate is mixed with potassium fluoride and produces potassium phosphate and magnesium fluoride. How many moles of each product is produced?



Calcium carbonate reacts with 14.8 g of hydrochloric acid. How many moles of calcium chloride are produced? Carbon dioxide and water are the other products.



How many grams of aluminum are needed to react with nitric acid (HNO3) in order to produce 0.129 g of hydrogen gas? Aluminum nitrate is the other product in the reaction.



Determine the number of moles of potassium bromide that need to react with chlorine gas in order to produce 6.4 moles of bromine. Potassium chloride is also a product in the reaction.



In the combustion of 16.2 moles of ethane (C2H6), how many moles of carbon dioxide are produced?


1
Expert's answer
2021-04-09T03:38:44-0400

PV = nRT

V = nRT / P

Mg(s)+2HCl(aq)→MgCl2(aq)+H2(g)

n(H2) = n (Mg) = m (Mg) / M (Mg) = 41.0 g / (24.0 g/mol) = 1.7 mol

R = 62360 mmHg/mol*K

T = 23 ˚C + 273 = 296 K

V(H2) = (1.7 mol ×62360 mmHg / mol*K * 296) / 850 mm Hg = 36917.1 ml = 36.9 L






C2H6 + 4 O2 → 2 C + 2 CO + 6 H2O

Moles of CO2= 2×16.2=32.4 moles